OpenStudy (anonymous):

14=7y-6x write the equation

OpenStudy (austinl):

That is an equation. I assume it is supposed to be in a particular form?

OpenStudy (austinl):

Hello....

OpenStudy (anonymous):

Welcome to openstudy friend but its an equation . What do you mean ?

OpenStudy (anonymous):

i agree with @austinL

OpenStudy (anonymous):

he is very smart

OpenStudy (anonymous):

kk,looks like its a mistaken quetsion

OpenStudy (anonymous):

They mean solve it for them. First find the variables cost, the number that the letters x and y stand for, until you've found the perfect couple in order to subtract it from one another enable to receive your answer of fourteen. Which should be fairly easy. x= 0 y= 2 therefore 14 = 7*2 - 6*0 7 * 2 = 14 6 * 0 = 0 14 - 0 = 14

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."